OpenStudy (anonymous):

Erik needs a plumber and must decide between two companies. For a service visit, Plumber A charges $85 to send a plumber, plus $20/h. Plumber B charges $60 to send a plumber, plus $30/h. The system that models this situation is given, where c is the cost of hiring a plumber and h is the number of hours the plumber works. The solution to the system is (2.5, 135). What is the correct interpretation of the solution of this system?

OpenStudy (anonymous):

c=85+20h c=60+30h A.Plumber A will charge more when he works 2.5 h, charging $135. B.Plumber B will charge more when she works 2.5 h, charging $135. C.Plumber A and Plumber B will both charge $135 if they each work 2.5 h. D.Plumber A and Plumber B will both charge $2.50 if they each work 135 h. please help will fan and medal